### v2

# Information geometry of Lévy processes and financial models

Abstract.
We explore the information geometry of Lévy processes. As a starting point, we derive the $\alpha$-divergence between two Lévy processes. Subsequently, the Fisher information matrix and the $\alpha$-connection associated with the geometry of Lévy processes are computed from the $\alpha$-divergence. In addition, we discuss statistical applications of this information geometry. As illustrative examples, we investigate the differential-geometric structures of various Lévy processes relevant to financial modeling, including tempered stable processes, the CGMY model, and variance gamma processes.
1. Introduction
Information geometry, which studies the differential-geometric structure of statistical models, has emerged as a powerful tool for understanding complex probabilistic systems. By representing families of probability distributions as Riemannian manifolds, this framework enables a geometric interpretation of statistical inference and parameter estimation procedures [15, 16, 17, 1, 25, 19]. The metric tensors and connections on statistical manifolds are derived from divergence functions, in particular, the $\alpha$-divergence, a generalization of the Kullback–Leibler divergence. In this context, the metric tensor corresponds to the Fisher information matrix, which plays a fundamental role in statistical theory and applications. In recent decades, information geometry has been successfully applied to exponential families [15, 16, 17, 26, 27, 1, 19], time series models [32, 31, 25, 35, 3, 4, 34, 12, 11, 29, 7], Bayesian prediction [25, 35, 34, 12, 11, 29], and, more recently, to stochastic processes with non-Gaussian features [8].
In previous work [8], the information geometry of tempered stable processes, a subclass of Lévy processes characterized by heavy tails and jumps, was investigated. The $\alpha$-divergence between two tempered stable processes was derived as a two-fold generalization of the result by Kim and Lee [21], in which the Kullback–Leibler divergence (relative entropy) for classical tempered stable processes, a subset of tempered stable processes, was obtained. Moreover, the corresponding Fisher information matrix and $\alpha$-connection, which capture the geometric structure of these processes, were constructed from the $\alpha$-divergence. Several statistical applications from the information geometry of tempered stable processes were also provided.
The differential-geometric approach to tempered stable processes opens new horizons for its applications in finance. Due to their ability to capture characteristic features of asset returns such as asymmetry and fat-tails, various tempered stable processes have been widely employed in financial modeling, as the normality assumption is empirically invalid in this domain. Their use in finance spans a range of models, including time series analysis via ARMA-GARCH models with tempered stable distributed innovations [22, 23, 24], as well as applications in portfolio management, risk management, and option pricing [36, 5, 10, 9, 18, 2, 20].
However, tempered stable processes, while rich and flexible for financial modeling, constitute only a subset of the broader class of Lévy processes. Lévy processes include not only various tempered stable processes but also stable, compound Poisson, variance gamma processes, and other models [14, 30, 28]. These processes are characterized by stationary and independent increments and serve as fundamental models in the study of jump processes and infinite divisibility. Extending the information geometry developed for tempered stable processes [8] to the full class of Lévy processes is a natural and necessary progression, with implications for both theoretical development and practical applications.
The present paper aims to undertake this extension. Our main goal is to develop the information geometry of Lévy processes. As mentioned above, we generalize the construction of information-geometric structures from the tempered stable process setting [8] to the full class of Lévy processes. This extension also provides an alternative generalization of the result in Cont and Tankov [14], where the Kullback–Leibler divergence for Lévy processes was derived. We begin by deriving the $\alpha$-divergence between two Lévy processes. Since divergence functions assess dissimilarity between probability distributions or statistical models, they underlie the definition of key geometric objects such as the Fisher information matrix and the $\alpha$-connection, derived from the $\alpha$-divergence [1, 13]. We also discuss potential statistical applications for Lévy processes from the geometric perspective.
Having established the general framework, we investigate several concrete examples of Lévy processes relevant to financial modeling. These include tempered stable processes [30], CGMY models [6], and variance gamma processes [28], all of which have been widely used to capture the stylized features of financial asset returns. For each model, we derive the $\alpha$-divergence and construct the associated information geometry such as Fisher information matrix and $\alpha$-connection. Moreover, we provide the statistical advantages originated from the geometry of each model.
By developing a unified geometric viewpoint on Lévy processes, this work contributes to the growing intersection of stochastic processes, mathematical finance, and statistics through the frame of information geometry. Our results offer both theoretical insights and practical tools for analyzing Lévy processes and jump-driven models. In doing so, we build a bridge between abstract geometric concepts and concrete applications, particularly in domains such as finance where Lévy processes play a central role.
This paper is organized as follows. In Section 2, we revisit the fundamentals of Lévy processes and review relevant prior work. Section 3 develops the information geometry of Lévy processes, including $\alpha$-divergence, Fisher information matrix, and $\alpha$-connection, and also discusses its statistical applications. In Section 4, we present several illustrative examples with financial applications. Finally, we conclude the paper with a summary.
2. Lévy processes
In this section, we provide the theoretical background on Lévy processes [33]. We also review prior work on the Kullback–Leibler divergence for Lévy processes [14]. These foundations will be extended in the next section to investigate the information geometry of Lévy processes.
Lévy processes constitute a broad and fundamental class of stochastic processes characterized by stationary and independent increments. By generalizing both Brownian motion and Poisson processes, they serve as canonical building blocks for modeling random phenomena that exhibit discontinuities, jumps, and heavy tails. This versatility makes Lévy processes particularly suitable for describing systems influenced by random shocks and bursts. They are especially important in fields such as quantitative finance, physics, signal processing, and biology, where non-Gaussian and discontinuous behavior is frequently observed.
Mathematically, a Lévy process $(X_{t},\mathbb{P})_{t\in[0,T]}$ on a probability space $(\Omega,\mathcal{F})$ is a stochastic process that satisfies the following properties:
  • $X_{0}=0$ almost surely.
  • Independent increments: For any $0\leq t_{0}\leq t_{1}\leq\cdots\leq t_{n}$, the increments $X_{t_{1}}-X_{t_{0}},X_{t_{2}}-X_{t_{1}},\cdots,X_{t_{n}}-X_{t_{n-1}}$ are mutually independent.
  • Stationary increments: For all $s,t\geq 0$, the distribution of $X_{t+s}-X_{s}$ depends only on $t$.
  • Stochastic continuity: For all $\epsilon>0$, $\lim_{h\to 0}\mathbb{P}(|X_{t+h}-X_{t}|>\epsilon)=0$.
These conditions ensure that Lévy processes are useful for modeling phenomena with both continuous and discontinuous sample paths. While Brownian motion accounts for continuous evolution, Lévy processes incorporate jumps via their Lévy measures, capturing more complex and realistic dynamics.
The Lévy–Itô decomposition states that a Lévy process can be uniquely expressed as the combination of a Brownian (diffusion) component, a drift term, and a pure jump component governed by a Lévy measure. This decomposition for a Lévy process is fully characterized by its characteristic (Lévy) triplet $(\sigma,\nu,\gamma)$, where $\sigma\geq 0$ is the diffusion coefficient, $\nu$ is a Lévy measure on $\mathbb{R}$, and $\gamma\in\mathbb{R}$ is the drift term. The Lévy measure $\nu$ is a positive measure satisfying:
(1) $\nu({0})=0,\quad\int_{-1}^{+1}x^{2}\nu(dx)<\infty,\quad\int_{|x|>1}\nu(dx)<\infty.$
A cornerstone of Lévy process theory is the Lévy–Khintchine formula, which characterizes Lévy processes via their characteristic functions [30]. According to the Lévy–Khintchine formula, the characteristic function of a Lévy process $(X_{t},\mathbb{P})_{t\in[0,T]}$ with Lévy triplet $(\sigma,\nu,\gamma)$ is given by
(2) $\phi_{t}(z;\boldsymbol{\xi})=\mathbb{E}[{\rm e}^{izX_{t}}]=\exp\left(-t\frac{\sigma^{2}z^{2}}{2}+it\gamma z+t\int_{-\infty}^{\infty}({\rm e}^{izx}-1-izx\mathbbm{1}_{|x|\leq 1})\nu(dx)\right),$
where $\boldsymbol{\xi}$ is the parameter vector of the Lévy process, and $\mathbbm{1}$ is the indicator function such that $\mathbbm{1}_{A}(x)=1$ when $x\in A$ and 0 otherwise.
The logarithmic characteristic function $\Phi_{t}(z;\boldsymbol{\xi})$, also known as the Lévy (or characteristic) exponent, is given by the Lévy–Khintchine representation:
(3) $\Phi_{t}(z;\boldsymbol{\xi})=\log\phi_{t}(z;\boldsymbol{\xi})=-t\frac{\sigma^{2}z^{2}}{2}+it\gamma z+t\int_{-\infty}^{\infty}({\rm e}^{izx}-1-izx\mathbbm{1}_{|x|\leq 1})\nu(dx).$
It is noteworthy that the martingale condition for Lévy processes is expressed as
(4) $\Phi_{t}(-i;\boldsymbol{\xi})=0,$
which is identically satisfied when
(5) $\gamma=-\frac{\sigma^{2}}{2}-\int_{-\infty}^{\infty}\left({\rm e}^{x}-1-x\mathbbm{1}_{|x|\leq 1}\right)\nu(dx).$
This martingale condition is particularly important in the context of exponential Lévy processes [14].
According to Sato [33], the Radon–Nikodym derivative between two Lévy processes can be defined under some conditions. The following proposition outlines the formulation of the Radon–Nikodym derivative for Lévy processes.
Proposition 1 (Sato (1999)).
Let $(X_{t},\mathbb{P})_{t\in[0,T]}$ and $(X_{t},\mathbb{Q})_{t\in[0,T]}$ be two Lévy processes on $(\Omega,\mathcal{F})$ with characteristic triplets $(\sigma_{\mathbb{P}},\nu_{\mathbb{P}},\gamma_{\mathbb{P}})$ and $(\sigma_{\mathbb{Q}},\nu_{\mathbb{Q}},\gamma_{\mathbb{Q}})$, respectively. Then $\mathbb{P}|_{\mathcal{F}_{t}}$ and $\mathbb{Q}|_{\mathcal{F}_{t}}$ are mutually absolutely continuous for all $t$ if and only if the following conditions are satisfied:
  • $\sigma_{\mathbb{P}}=\sigma_{\mathbb{Q}}=\sigma$.
  • The Lévy measures are mutually absolutely continuous with
    (6) $\int_{-\infty}^{\infty}({\rm e}^{\psi(x)/2}-1)^{2}\nu_{\mathbb{Q}}(dx)<\infty,$
    where $\psi$ is the logarithm of the Radon–Nikodym density of $\nu_{\mathbb{P}}$ with respect to $\nu_{\mathbb{Q}}$: ${\rm e}^{\psi}(x)=d\nu_{\mathbb{P}}/d\nu_{\mathbb{Q}}$.
  • If $\sigma=0$, then in addition, $\gamma_{\mathbb{Q}}$ must satisfy
    (7) $\gamma_{\mathbb{P}}-\gamma_{\mathbb{Q}}=\int_{-1}^{1}x(\nu_{\mathbb{P}}-\nu_{\mathbb{Q}})(dx).$
The Radon–Nikodym derivative is given by
(8) $\frac{d\mathbb{P}|_{\mathcal{F}_{t}}}{d\mathbb{Q}|_{\mathcal{F}_{t}}}={\rm e}^{U_{t}},$
where $(U_{t},\mathbb{U})_{t\in[0,T]}$ is a Lévy process with characteristic triplet $(\sigma_{\mathbb{U}},\nu_{\mathbb{U}},\gamma_{\mathbb{U}})$ given by
(9) $\sigma_{\mathbb{U}}$ $=\sigma\eta,$
(10) $\nu_{\mathbb{U}}$ $=\nu_{\mathbb{Q}}\psi^{-1},$
(11) $\gamma_{\mathbb{U}}$ $=-\frac{1}{2}(\sigma\eta)^{2}-\int_{-\infty}^{\infty}({\rm e}^{y}-1-y\mathbbm{1}_{|y|\leq 1})(\nu_{\mathbb{Q}}\psi^{-1})(dy),$
and $\eta$ is chosen so that
(12) $\gamma_{\mathbb{P}}-\gamma_{\mathbb{Q}}-\int_{-1}^{1}x(\nu_{\mathbb{P}}-\nu_{\mathbb{Q}})(dx)=\sigma^{2}\eta.$
Moreover, $U_{t}$ satisfies
(13) $\mathbb{E}_{\mathbb{Q}}[{\rm e}^{U_{t}}]=1$
for all $t$.
We assume that Lévy processes in this paper satisfy Proposition 1.
Given the Radon–Nikodym derivative and the conditions stated in Proposition 1, the Kullback–Leibler divergence between Lévy processes can be well-defined. Cont and Tankov [14] derived an explicit expression for the Kullback–Leibler divergence of Lévy processes in this context.
Proposition 2 (Cont and Tankov (2004)).
Let $(X_{t},\mathbb{P})_{t\in[0,T]}$ and $(X_{t},\mathbb{Q})_{t\in[0,T]}$ be two Lévy processes on $(\Omega,\mathcal{F})$ with characteristic triplets $(\sigma,\nu_{\mathbb{P}},\gamma_{\mathbb{P}})$ and $(\sigma,\nu_{\mathbb{Q}},\gamma_{\mathbb{Q}})$. The Kullback–Leibler divergence $KL(\mathbb{P}||\mathbb{Q})$ is then given by
(14) $KL(\mathbb{P}||\mathbb{Q})=\frac{T}{2\sigma^{2}}\big(\gamma_{\mathbb{P}}-\gamma_{\mathbb{Q}}-\int_{-1}^{1}x(\nu_{\mathbb{P}}-\nu_{\mathbb{Q}})(dx)\big)^{2}+T\int_{-\infty}^{\infty}\Big(\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)\log{\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+1\Big)\nu_{\mathbb{Q}}(dx).$
If $\mathbb{P}$ and $\mathbb{Q}$ correspond to risk-neutral exponential Lévy models, i.e., verify the martingale condition given by Eq. (5), the Kullback–Leibler divergence is reduced to
(15) $KL(\mathbb{P}||\mathbb{Q})=\frac{T}{2\sigma^{2}}\big(\int_{-\infty}^{\infty}({\rm e}^{x}-1)(\nu_{\mathbb{P}}-\nu_{\mathbb{Q}})(dx)\big)^{2}+T\int_{-\infty}^{\infty}\Big(\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)\log{\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+1\Big)\nu_{\mathbb{Q}}(dx).$
In brief, the proof relies on applying the Radon–Nikodym derivative from Proposition 1 to the definition of the Kullback–Leibler divergence. We will revisit this proof in the following section.
3. Information geometry of Lévy processes
In this section, we explore the information geometry of Lévy processes. We begin by deriving the $\alpha$-divergence of Lévy processes from the definition of $f$-divergence. We then investigate the associated geometric structures, including Fisher information matrix and $\alpha$-connection, and discuss statistical applications.
[…]
where $f$ is a convex function such that $f(1)=0$ and $f(x)$ is finite for all $x>0$.
The $f$-divergence covers a variety of distance and divergence measures commonly used in statistics and information theory. For example, the Kullback–Leibler divergence arises from the $f$-divergence when the function $f$ is given by
(17) $f(t)=t\log{t}-t+1.$
Another important example of $f$-divergence is the $\alpha$-divergence, which plays a central role in information geometry [1]. The $\alpha$-divergence can be expressed as the $f$-divergence corresponding to the following choice of $f$:
(21) $f^{(\alpha)}(t)=\left\{\begin{array}[]{ll}\frac{4}{1-\alpha^{2}}\Big(\frac{1-\alpha}{2}t+\frac{1+\alpha}{2}-t^{\frac{1-\alpha}{2}}\Big)&(\alpha\neq\pm 1)\\ t\log{t}-t+1&(\alpha=-1)\\ -\log{t}+t-1&(\alpha=1)\end{array}\right.$.
It is straightforward to verify that the functions $f^{(\alpha)}$ for $\alpha=\pm 1$ can be obtained by applying L’Hôpital’s rule to the $f$-divergence for $\alpha\neq\pm 1$ and taking the limit as $\alpha\to\pm 1$, respectively. Moreover, the $\alpha=-1$ case in Eq. (21) coincides with Eq. (17), therefore recovering the Kullback–Leibler divergence.
[…]
(25) $D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=\left\{\begin{array}[]{ll}\frac{4}{1-\alpha^{2}}\int_{-\infty}^{\infty}(\frac{1-\alpha}{2}p(x;\boldsymbol{\xi})+\frac{1+\alpha}{2}q(x;\tilde{\boldsymbol{\xi}})-p(x;\boldsymbol{\xi})^{\frac{1-\alpha}{2}}q(x;\tilde{\boldsymbol{\xi}})^{\frac{1+\alpha}{2}})dx&(\alpha\neq\pm 1)\\ \int_{-\infty}^{\infty}(p(x;\boldsymbol{\xi})\log{\frac{p(x;\boldsymbol{\xi})}{q(x;\tilde{\boldsymbol{\xi}})}}-p(x;\boldsymbol{\xi})+q(x;\tilde{\boldsymbol{\xi}}))dx&(\alpha=-1)\\ \int_{-\infty}^{\infty}(q(x;\tilde{\boldsymbol{\xi}})\log{\frac{q(x;\tilde{\boldsymbol{\xi}})}{p(x;\boldsymbol{\xi})}}-q(x;\tilde{\boldsymbol{\xi}})+p(x;\boldsymbol{\xi}))dx&(\alpha=1)\end{array}\right$.,
where $\boldsymbol{\xi}$ and $\tilde{\boldsymbol{\xi}}$ are parameters of $p$ and $q$, respectively.
It is noteworthy that the $\alpha$-divergence exhibits an interesting property known as $\alpha$-duality [1]:
(26) $D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=D^{(-\alpha)}(\mathbb{Q}||\mathbb{P}).$
Flipping the sign of $\alpha$ is equivalent to exchanging the roles of $\mathbb{P}$ and $\mathbb{Q}$. In particular, it is straightforward to verify that the Hellinger distance, corresponding to $\alpha=0$, is self-dual.
It is natural to extend the preceding discussion from probability density functions to Radon–Nikodym derivatives. The $f$-divergence can be defined in terms of the Radon–Nikodym derivative as follows:
(27) $D_{f}(\mathbb{P}||\mathbb{Q})=\mathbb{E}_{\mathbb{Q}}\Big[f\Big(\frac{d\mathbb{P}}{d\mathbb{Q}}\Big)\Big].$
Based on the definition of the Kullback–Leibler divergence given by Eq. (27) together with Eq. (17), Cont and Tankov [14] computed the Kullback–Leibler divergence for Lévy processes under Proposition 1 by evaluating the following expression:
(28) $KL(\mathbb{P}||\mathbb{Q})=\mathbb{E}_{\mathbb{Q}}[f({\rm{e}}^{U_{T}})],$
where $f$ is given by Eq. (17) and $U_{T}$ is defined in Proposition 1. This calculation consists of the basis for the proof of Proposition 2, as mentioned earlier.
Analogous to Proposition 2, the $\alpha$-divergence between Lévy processes can be derived by applying Eq. (21) to Eq. (27):
(29) $D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=\mathbb{E}_{\mathbb{Q}}[f^{(\alpha)}({\rm{e}}^{U_{T}})].$
Before proceeding with the derivation of the $\alpha$-divergence, we first consider the following proposition for further discussion.
Proposition 3.
[…]
(30) $\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})=-\Phi_{T}(-i\frac{1-\alpha}{2};\boldsymbol{\xi}_{\mathbb{U}}),$
where $\Phi_{T}(z;\boldsymbol{\xi})$ is given by Eq. (3) and $U_{T}$ is defined in Proposition 1.
[…]
(31) $\begin{aligned} \Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})=&\frac{1-\alpha^{2}}{4}\frac{T}{2\sigma^{2}}\big(\gamma_{\mathbb{P}}-\gamma_{\mathbb{Q}}-\int_{-1}^{1}x(\nu_{\mathbb{P}}-\nu_{\mathbb{Q}})(dx)\big)^{2}\\ &+T\int_{-\infty}^{\infty}\Big(\frac{1-\alpha}{2}\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+\frac{1+\alpha}{2}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)^{\frac{1-\alpha}{2}}\Big)\nu_{\mathbb{Q}}(dx).\end{aligned}$
[…]
(32) $\begin{aligned} \Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})=&\frac{1-\alpha^{2}}{4}\frac{T}{2\sigma^{2}}\big(\int_{-\infty}^{\infty}({\rm e}^{x}-1)(\nu_{\mathbb{P}}-\nu_{\mathbb{Q}})(dx)\big)^{2}\\ &+T\int_{-\infty}^{\infty}\Big(\frac{1-\alpha}{2}\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+\frac{1+\alpha}{2}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)^{\frac{1-\alpha}{2}}\Big)\nu_{\mathbb{Q}}(dx).\end{aligned}$
When $\sigma=0$, we have the $\alpha$-divergence between two Lévy processes as
(33) $\begin{aligned} \Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})=T\int_{-\infty}^{\infty}\Big(\frac{1-\alpha}{2}\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+\frac{1+\alpha}{2}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)^{\frac{1-\alpha}{2}}\Big)\nu_{\mathbb{Q}}(dx).\end{aligned}$
Proof.
When $\sigma\neq 0$, $\Phi_{T}(-i\frac{1-\alpha}{2};\boldsymbol{\xi}_{\mathbb{U}})$ is obtained as
$T^{-1}\Phi_{T}(-i\frac{1-\alpha}{2};\boldsymbol{\xi}_{\mathbb{U}})=\Big(\frac{1-\alpha}{2}\Big)^{2}\frac{\sigma_{\mathbb{U}}^{2}}{2}+\frac{1-\alpha}{2}\gamma_{\mathbb{U}}+\int_{-\infty}^{\infty}({\rm e}^{\frac{1-\alpha}{2}y}-1-\frac{1-\alpha}{2}y\mathbbm{1}_{|y|\leq 1})\nu_{\mathbb{U}}(dy).$
By applying Eq. (9), Eq. (10), Eq. (11), and Eq. (12), $T^{-1}\Phi_{T}\left(-i\frac{1-\alpha}{2};\boldsymbol{\xi}_{\mathbb{U}}\right)$ can be expressed as
$T^{-1}\Phi_{T}(-i\frac{1-\alpha}{2};\boldsymbol{\xi}_{\mathbb{U}})$
$=\Big(\frac{1-\alpha}{2}\Big)^{2}\frac{(\sigma\eta)^{2}}{2}+\frac{1-\alpha}{2}\Big(-\frac{(\sigma\eta)^{2}}{2}-\int_{-\infty}^{\infty}({\rm e}^{y}-1-y\mathbbm{1}_{|y|\leq 1})(\nu_{\mathbb{Q}}\psi^{-1})(dy)\Big)$
 $\quad\textrm{ }+\int_{-\infty}^{\infty}({\rm e}^{\frac{1-\alpha}{2}y}-1-\frac{1-\alpha}{2}y\mathbbm{1}_{|y|\leq 1})(\nu_{\mathbb{Q}}\psi^{-1})(dy)$
$=-\frac{1-\alpha^{2}}{4}\frac{(\sigma\eta)^{2}}{2}-\int_{-\infty}^{\infty}(\frac{1-\alpha}{2}{\rm e}^{y}+\frac{1+\alpha}{2}-{\rm e}^{\frac{1-\alpha}{2}y})(\nu_{\mathbb{Q}}\psi^{-1})(dy)$
$=-\frac{1-\alpha^{2}}{4}\frac{(\sigma\eta)^{2}}{2}-\int_{-\infty}^{\infty}\Big(\frac{1-\alpha}{2}\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+\frac{1+\alpha}{2}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)^{\frac{1-\alpha}{2}}\Big)\nu_{\mathbb{Q}}(dx)$
$=-\frac{1-\alpha^{2}}{4}\frac{1}{2\sigma^{2}}\big(\gamma_{\mathbb{P}}-\gamma_{\mathbb{Q}}-\int_{-1}^{1}x(\nu_{\mathbb{P}}-\nu_{\mathbb{Q}})(dx)\big)^{2}-\int_{-\infty}^{\infty}\Big(\frac{1-\alpha}{2}\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+\frac{1+\alpha}{2}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)^{\frac{1-\alpha}{2}}\Big)\nu_{\mathbb{Q}}(dx).$
From the equation above, we derive Eq. (31).
If $\mathbb{P}$ and $\mathbb{Q}$ are risk-neutral exponential Lévy models, the martingale condition in Eq. (5) is satisfied. By applying this condition to the first term in Eq. (31), the expression simplifies to
$\gamma_{\mathbb{P}}-\gamma_{\mathbb{Q}}-\int_{-1}^{1}x(\nu_{\mathbb{P}}-\nu_{\mathbb{Q}})(dx)$
$=\big(-\frac{\sigma^{2}}{2}-\int_{-\infty}^{\infty}({\rm e}^{x}-1-x\mathbbm{1}_{|x|\leq 1})\nu_{\mathbb{P}}(dx)\big)-\big(-\frac{\sigma^{2}}{2}-\int_{-\infty}^{\infty}({\rm e}^{x}-1-x\mathbbm{1}_{|x|\leq 1})\nu_{\mathbb{Q}}(dx)\big)-\int_{-1}^{1}x(\nu_{\mathbb{P}}-\nu_{\mathbb{Q}})(dx)$
$=-\int_{-\infty}^{\infty}({\rm e}^{x}-1)(\nu_{\mathbb{P}}-\nu_{\mathbb{Q}})(dx).$
Using this term, it is straightforward to verify that substituting the expression above into Eq. (31) yields Eq. (32).
When $\sigma=0$, $\Phi_{T}(-i\frac{1-\alpha}{2};\boldsymbol{\xi}_{\mathbb{U}})$ is obtained as
$T^{-1}\Phi_{T}(-i\frac{1-\alpha}{2};\boldsymbol{\xi}_{\mathbb{U}})$ $=\int_{-\infty}^{\infty}({\rm e}^{\frac{1-\alpha}{2}y}-1-\frac{1-\alpha}{2}y\mathbbm{1}_{|y|\leq 1})\nu_{\mathbb{U}}(dy)$
$=-\int_{-\infty}^{\infty}\Big(\frac{1-\alpha}{2}\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+\frac{1+\alpha}{2}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)^{\frac{1-\alpha}{2}}\Big)\nu_{\mathbb{Q}}(dx).$
From the equation above, we derive Eq. (33). This also implies that when $\sigma=0$, the first terms in both Eq. (31) and Eq. (32) vanish, allowing Eq. (33) to be obtained directly from them. ∎
[…]
The following theorem provides the expression for the $\alpha$-divergence of Lévy processes.
Theorem 1.
Let $(X_{t},\mathbb{P})_{t\in[0,T]}$ and $(X_{t},\mathbb{Q})_{t\in[0,T]}$ be Lévy processes on $(\Omega,\mathcal{F})$ with Lévy triplets of $(\sigma,\nu_{\mathbb{P}},\gamma_{\mathbb{P}})$ and $(\sigma,\nu_{\mathbb{Q}},\gamma_{\mathbb{Q}})$, respectively. When $\sigma\neq 0$, we have the $\alpha$-divergence between two Lévy processes as
(40) $D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=\left\{\begin{array}[]{ll}\frac{4}{1-\alpha^{2}}\Big\{1-\exp\big[-\frac{1-\alpha^{2}}{4}\frac{T}{2\sigma^{2}}\big(\gamma_{\mathbb{P}}-\gamma_{\mathbb{Q}}-\int_{-1}^{1}x(\nu_{\mathbb{P}}-\nu_{\mathbb{Q}})(dx)\big)^{2}&\\ -T\int_{-\infty}^{\infty}\Big(\frac{1-\alpha}{2}\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+\frac{1+\alpha}{2}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)^{\frac{1-\alpha}{2}}\Big)\nu_{\mathbb{Q}}(dx)\big]\Big\}&(\alpha\neq\pm 1)\\ \frac{T}{2\sigma^{2}}\big(\gamma_{\mathbb{P}}-\gamma_{\mathbb{Q}}-\int_{-1}^{1}x(\nu_{\mathbb{P}}-\nu_{\mathbb{Q}})(dx)\big)^{2}&\\ +T\int_{-\infty}^{\infty}\Big(\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)\log{\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+1\Big)\nu_{\mathbb{Q}}(dx)&(\alpha=-1)\\ \frac{T}{2\sigma^{2}}\big(\gamma_{\mathbb{Q}}-\gamma_{\mathbb{P}}-\int_{-1}^{1}x(\nu_{\mathbb{Q}}-\nu_{\mathbb{P}})(dx)\big)^{2}&\\ +T\int_{-\infty}^{\infty}\Big(\Big(\frac{d\nu_{\mathbb{Q}}}{d\nu_{\mathbb{P}}}\Big)\log{\Big(\frac{d\nu_{\mathbb{Q}}}{d\nu_{\mathbb{P}}}\Big)}-\Big(\frac{d\nu_{\mathbb{Q}}}{d\nu_{\mathbb{P}}}\Big)+1\Big)\nu_{\mathbb{P}}(dx)&(\alpha=1)\end{array}\right.$.
If $\mathbb{P}$ and $\mathbb{Q}$ correspond to risk-neutral exponential Lévy models, i.e., verify the martingale condition given by Eq. (5), the $\alpha$-divergence between two Lévy processes is represented with
(47) $D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=\left\{\begin{array}[]{ll}\frac{4}{1-\alpha^{2}}\Big\{1-\exp\big[-\frac{1-\alpha^{2}}{4}\frac{T}{2\sigma^{2}}\big(\int_{-\infty}^{\infty}({\rm e}^{x}-1)(\nu_{\mathbb{P}}-\nu_{\mathbb{Q}})(dx)\big)^{2}&\\ -T\int_{-\infty}^{\infty}\Big(\frac{1-\alpha}{2}\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+\frac{1+\alpha}{2}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)^{\frac{1-\alpha}{2}}\Big)\nu_{\mathbb{Q}}(dx)\big]\Big\}&(\alpha\neq\pm 1)\\ \frac{T}{2\sigma^{2}}\big(\int_{-\infty}^{\infty}({\rm e}^{x}-1)(\nu_{\mathbb{P}}-\nu_{\mathbb{Q}})(dx)\big)^{2}&\\ +T\int_{-\infty}^{\infty}\Big(\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)\log{\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+1\Big)\nu_{\mathbb{Q}}(dx)&(\alpha=-1)\\ \frac{T}{2\sigma^{2}}\big(\int_{-\infty}^{\infty}({\rm e}^{x}-1)(\nu_{\mathbb{Q}}-\nu_{\mathbb{P}})(dx)\big)^{2}&\\ +T\int_{-\infty}^{\infty}\Big(\Big(\frac{d\nu_{\mathbb{Q}}}{d\nu_{\mathbb{P}}}\Big)\log{\Big(\frac{d\nu_{\mathbb{Q}}}{d\nu_{\mathbb{P}}}\Big)}-\Big(\frac{d\nu_{\mathbb{Q}}}{d\nu_{\mathbb{P}}}\Big)+1\Big)\nu_{\mathbb{P}}(dx)&(\alpha=1)\end{array}\right.$.
When $\sigma=0$, we have the $\alpha$-divergence between two Lévy processes as
(51) $D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=\left\{\begin{array}[]{ll}\frac{4}{1-\alpha^{2}}\Big\{1-\exp\big[-T\int_{-\infty}^{\infty}\Big(\frac{1-\alpha}{2}\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+\frac{1+\alpha}{2}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)^{\frac{1-\alpha}{2}}\Big)\nu_{\mathbb{Q}}(dx)\big]\Big\}&(\alpha\neq\pm 1)\\ T\int_{-\infty}^{\infty}\Big(\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)\log{\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+1\Big)\nu_{\mathbb{Q}}(dx)&(\alpha=-1)\\ T\int_{-\infty}^{\infty}\Big(\Big(\frac{d\nu_{\mathbb{Q}}}{d\nu_{\mathbb{P}}}\Big)\log{\Big(\frac{d\nu_{\mathbb{Q}}}{d\nu_{\mathbb{P}}}\Big)}-\Big(\frac{d\nu_{\mathbb{Q}}}{d\nu_{\mathbb{P}}}\Big)+1\Big)\nu_{\mathbb{P}}(dx)&(\alpha=1)\end{array}\right.$.
Proof.
We consider the cases $\sigma\neq 0$ and $\sigma=0$ separately.
We begin the proof with the case $\sigma\neq 0$. First, the $\alpha$-divergence for $\alpha\neq\pm 1$ is derived by evaluating Eq. (29):
$D^{(\alpha)}(\mathbb{P}||\mathbb{Q})$ $=\frac{4}{1-\alpha^{2}}\mathbb{E}_{\mathbb{Q}}[\frac{1-\alpha}{2}{\rm{e}}^{U_{T}}+\frac{1+\alpha}{2}-{\rm{e}}^{\frac{1-\alpha}{2}U_{T}}].$
By using the logarithmic characteristic function $\Phi_{t}(z;\boldsymbol{\xi})$ from Eq. (3), the $\alpha$-divergence for $\alpha\neq\pm 1$ can be expressed as
(52) $D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=\frac{4}{1-\alpha^{2}}\Big(1-\mathbb{E}_{\mathbb{Q}}[{\rm{e}}^{\frac{1-\alpha}{2}U_{T}}]\Big)=\frac{4}{1-\alpha^{2}}\Big(1-{\rm{e}}^{\Phi_{T}(-i\frac{1-\alpha}{2};\boldsymbol{\xi}_{\mathbb{U}})}\Big).$
[…]
By substituting Eq. (31) into Eq. (53), the $\alpha$-divergence for $\alpha\neq\pm 1$ is given by
$D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=\frac{4}{1-\alpha^{2}}\Big\{1$- $\exp\big[-\frac{1-\alpha^{2}}{4}\frac{T}{2\sigma^{2}}\big(\gamma_{\mathbb{P}}-\gamma_{\mathbb{Q}}-\int_{-1}^{1}x(\nu_{\mathbb{P}}-\nu_{\mathbb{Q}})(dx)\big)^{2}$
$-T\int_{-\infty}^{\infty}\Big(\frac{1-\alpha}{2}\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+\frac{1+\alpha}{2}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)^{\frac{1-\alpha}{2}}\Big)\nu_{\mathbb{Q}}(dx)\big]\Big\}.$
When $\alpha=-1$, the Kullback–Leibler divergence can be obtained in two different ways. The first approach is a direct calculation by plugging the $\alpha=-1$ case in Eq. (21) into Eq. (29), which yields the same result with the derivation presented by Cont and Tankov [14]. The second method involves applying L’Hôpital’s rule to the expression for $\alpha$-divergence with $\alpha\neq\pm 1$ and taking the limit as $\alpha\to-1$, which also recovers the Kullback–Leibler divergence.
We begin with the first solution, given by:
$D^{(-1)}(\mathbb{P}||\mathbb{Q})=\mathbb{E}_{\mathbb{Q}}[U_{T}{\rm{e}}^{U_{T}}-{\rm{e}}^{U_{T}}+1]=\mathbb{E}_{\mathbb{Q}}[U_{T}{\rm{e}}^{U_{T}}]$
$=-i\frac{d}{dz}\mathbb{E}_{\mathbb{Q}}[{\rm{e}}^{izU_{T}}]|_{z=-i}=-i\Phi^{\prime}_{T}(-i;\boldsymbol{\xi}_{\mathbb{U}}){\rm{e}}^{\Phi_{T}(-i;\boldsymbol{\xi}_{\mathbb{U}})}=-i\Phi^{\prime}_{T}(-i;\boldsymbol{\xi}_{\mathbb{U}}),$
where $\Phi^{\prime}_{T}(z;\boldsymbol{\xi}_{\mathbb{U}})$ is directly obtained by differentiating Eq. (3):
$\Phi^{\prime}_{T}(z;\boldsymbol{\xi}_{\mathbb{U}})=-T\sigma^{2}_{\mathbb{U}}z+iT\gamma_{\mathbb{U}}+T\int_{-\infty}^{\infty}(iy{\rm e}^{izy}-iy\mathbbm{1}_{|y|\leq 1})\nu_{\mathbb{U}}(dy).$
By plugging Eq. (9), Eq. (10), Eq. (11), and Eq. (12), $\Phi^{\prime}_{T}(-i;\boldsymbol{\xi}_{\mathbb{U}})$ is obtained as
$-iT^{-1}\Phi^{\prime}_{T}(-i;\boldsymbol{\xi}_{\mathbb{U}})=\sigma^{2}_{\mathbb{U}}+\gamma_{\mathbb{U}}+\int_{-\infty}^{\infty}(y{\rm e}^{y}-y\mathbbm{1}_{|y|\leq 1})\nu_{\mathbb{U}}(dy)$
$=(\sigma\eta)^{2}+\Big(-\frac{(\sigma\eta)^{2}}{2}-\int_{-\infty}^{\infty}({\rm e}^{y}-1-y\mathbbm{1}_{|y|\leq 1})(\nu_{\mathbb{Q}}\psi^{-1})(dy)\Big)+\int_{-\infty}^{\infty}(y{\rm e}^{y}-y\mathbbm{1}_{|y|\leq 1})(\nu_{\mathbb{Q}}\psi^{-1})(dy)\Big)$
$=\frac{(\sigma\eta)^{2}}{2}+\int_{-\infty}^{\infty}(y{\rm e}^{y}-{\rm e}^{y}+1)(\nu_{\mathbb{Q}}\psi^{-1})(dy)$
$=\frac{(\sigma\eta)^{2}}{2}+\int_{-\infty}^{\infty}\Big(\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)\log{\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)}+1-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)\Big)\nu_{\mathbb{Q}}(dx)$
$=\frac{1}{2\sigma^{2}}\big(\gamma_{\mathbb{P}}-\gamma_{\mathbb{Q}}-\int_{-1}^{1}x(\nu_{\mathbb{P}}-\nu_{\mathbb{Q}})(dx)\big)^{2}+\int_{-\infty}^{\infty}\Big(\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)\log{\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)}+1-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)\Big)\nu_{\mathbb{Q}}(dx).$
From the calculation above, the Kullback–Leibler divergence can be expressed as
$D^{(-1)}(\mathbb{P}||\mathbb{Q})=\frac{T}{2\sigma^{2}}\big(\gamma_{\mathbb{P}}-\gamma_{\mathbb{Q}}-\int_{-1}^{1}x(\nu_{\mathbb{P}}-\nu_{\mathbb{Q}})(dx)\big)^{2}+T\int_{-\infty}^{\infty}\Big(\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)\log{\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)}+1-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)\Big)\nu_{\mathbb{Q}}(dx).$
The same result can also be obtained by applying L’Hôpital’s rule.
Similarly, there are two ways to obtain the $1$-divergence: one by leveraging L’Hôpital’s rule, and the other by direct calculation from the definition of $\alpha$-divergence. Additionally, a third approach involves using $\alpha$-duality to derive the dual Kullback–Leibler divergence:
$D^{(1)}(\mathbb{P}||\mathbb{Q})=D^{(-1)}(\mathbb{Q}||\mathbb{P})$
$=\frac{T}{2\sigma^{2}}\big(\gamma_{\mathbb{Q}}-\gamma_{\mathbb{P}}-\int_{-1}^{1}x(\nu_{\mathbb{Q}}-\nu_{\mathbb{P}})(dx)\big)^{2}+T\int_{-\infty}^{\infty}\Big(\Big(\frac{d\nu_{\mathbb{Q}}}{d\nu_{\mathbb{P}}}\Big)\log{\Big(\frac{d\nu_{\mathbb{Q}}}{d\nu_{\mathbb{P}}}\Big)}+1-\Big(\frac{d\nu_{\mathbb{Q}}}{d\nu_{\mathbb{P}}}\Big)\Big)\nu_{\mathbb{P}}(dx).$
It is straightforward to verify that the same result is obtained from the other two methods: applying L’Hôpital’s rule and performing a direct calculation based on the definition of $\alpha$-divergence.
We now return to the case of $\sigma=0$. As noted earlier, when $\sigma=0$, the first terms in Eq. (31) and Eq. (32) vanish by Proposition 3, allowing Eq. (51) to be derived directly from Eq. (40) and Eq. (47):
$D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=\frac{4}{1-\alpha^{2}}\Big\{1-\exp\big[-T\int_{-\infty}^{\infty}\Big(\frac{1-\alpha}{2}\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+\frac{1+\alpha}{2}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)^{\frac{1-\alpha}{2}}\Big)\nu_{\mathbb{Q}}(dx)\big]\Big\}.$
Similarly, the Kullback–Leibler divergence can be obtained as
$D^{(-1)}(\mathbb{P}||\mathbb{Q})=T\int_{-\infty}^{\infty}\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\log{\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)}+1-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)\Big)\nu_{\mathbb{Q}}(dx),$
and the corresponding dual divergence is given by
$D^{(1)}(\mathbb{P}||\mathbb{Q})=T\int_{-\infty}^{\infty}\Big(\frac{d\nu_{\mathbb{Q}}}{d\nu_{\mathbb{P}}}\log{\Big(\frac{d\nu_{\mathbb{Q}}}{d\nu_{\mathbb{P}}}\Big)}+1-\Big(\frac{d\nu_{\mathbb{Q}}}{d\nu_{\mathbb{P}}}\Big)\Big)\nu_{\mathbb{P}}(dx).$
It is worth noting that the Kullback–Leibler divergences in Eq. (40) and Eq. (47), i.e., the cases with $\alpha=-1$, are identical to those presented in Proposition 2, as derived by Cont and Tankov [14].
[…]
Corollary 1.
[…]
(57) $D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=\left\{\begin{array}[]{ll}\frac{4}{1-\alpha^{2}}\Big(1-{\rm{e}}^{-\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})}\Big).&(\alpha\neq\pm 1)\\ \lim_{\alpha\to-1}-\frac{2}{\alpha}\frac{\partial_{\alpha}\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})}{\partial\alpha}&(\alpha=-1)\\ \lim_{\alpha\to 1}-\frac{2}{\alpha}\frac{\partial_{\alpha}\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})}{\partial\alpha}&(\alpha=1)\end{array}\right.$.
Proof.
The proof is straightforward. For $\alpha\neq\pm 1$, the result follows directly from Eq. (53). When $\alpha=\pm 1$, as discussed in the proof of Theorem 1, we apply L’Hôpital’s rule to Eq. (53) and take the limit as $\alpha\to\pm 1$:
(58) $D^{(\pm 1)}(\mathbb{P}||\mathbb{Q})=\lim_{\alpha\to\pm 1}-\frac{2}{\alpha}\frac{\partial_{\alpha}\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})}{\partial\alpha}{\rm e}^{-\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})}$.
[…]
This approximation also yields the same values of the $\alpha$-divergence in the limits as $\alpha\to\pm 1$.
Having obtained the $\alpha$-divergence for Lévy processes, we can now derive associated information-geometric structures, such as Fisher information matrix and $\alpha$-connection, from the $\alpha$-divergence.
Theorem 2.
Let $(X_{t},\mathbb{P})_{t\in[0,T]}$ be a Lévy process on $(\Omega,\mathcal{F})$ with the Lévy triplet of $(\sigma,\nu,\gamma)$. When $\sigma\neq 0$, the Fisher information matrix of Lévy processes is represented with
(60) $g_{ij}=\frac{T}{\sigma^{2}}\partial_{i}\big(\gamma-\int_{-1}^{1}x\nu(dx)\big)\partial_{j}\big(\gamma-\int_{-1}^{1}x\nu(dx)\big)+T\int_{-\infty}^{\infty}\partial_{i}\log{\Big(\frac{d\nu}{dx}\Big)}\partial_{j}\log{\Big(\frac{d\nu}{dx}\Big)}\nu(dx),$
and the $\alpha$-connection for information geometry of Lévy processes is given as
(61) $\begin{split}\Gamma^{(\alpha)}_{ij,k}=&\frac{T}{\sigma^{2}}\partial_{i}\partial_{j}\big(\gamma-\int_{-1}^{1}x\nu(dx)\big)\partial_{k}\big(\gamma-\int_{-1}^{1}x\nu(dx)\big)\\ &+T\int_{-\infty}^{\infty}\Big(\partial_{i}\partial_{j}\log{\Big(\frac{d\nu}{dx}\Big)}+\frac{1-\alpha}{2}\partial_{i}\log{\Big(\frac{d\nu}{dx}\Big)}\partial_{j}\log{\Big(\frac{d\nu}{dx}\Big)}\Big)\partial_{k}\log{\Big(\frac{d\nu}{dx}\Big)}\nu(dx),\end{split}$
where $i,j,$ and $k$ run for the coordinate system $\boldsymbol{\xi}$.
If $\mathbb{P}$ and $\mathbb{Q}$ correspond to risk-neutral exponential Lévy models, i.e., verify the martingale condition given by Eq. (5), the Fisher information matrix of Lévy processes is represented with
(62) $g_{ij}=$ $\frac{T}{\sigma^{2}}\partial_{i}\big(\int_{-\infty}^{\infty}({\rm e}^{x}-1)\nu(dx)\big)\partial_{j}\big(\int_{-\infty}^{\infty}({\rm e}^{x}-1)\nu(dx)\big)+T\int_{-\infty}^{\infty}\partial_{i}\log{\Big(\frac{d\nu}{dx}\Big)}\partial_{j}\log{\Big(\frac{d\nu}{dx}\Big)}\nu(dx),$
and the $\alpha$-connection for information geometry of Lévy processes is given as
(63) $\begin{split}\Gamma^{(\alpha)}_{ij,k}=&\frac{T}{\sigma^{2}}\partial_{i}\partial_{j}\big(\int_{-\infty}^{\infty}({\rm e}^{x}-1)\nu(dx)\big)\partial_{k}\big(\int_{-\infty}^{\infty}({\rm e}^{x}-1)\nu(dx)\big)\\ &+T\int_{-\infty}^{\infty}\Big(\partial_{i}\partial_{j}\log{\Big(\frac{d\nu}{dx}\Big)}+\frac{1-\alpha}{2}\partial_{i}\log{\Big(\frac{d\nu}{dx}\Big)}\partial_{j}\log{\Big(\frac{d\nu}{dx}\Big)}\Big)\partial_{k}\log{\Big(\frac{d\nu}{dx}\Big)}\nu(dx).\end{split}$
When $\sigma=0$, the Fisher information matrix and the $\alpha$-connection for information geometry of Lévy processes are represented with
(64) $g_{ij}=$ $T\int_{-\infty}^{\infty}\partial_{i}\log{\Big(\frac{d\nu}{dx}\Big)}\partial_{j}\log{\Big(\frac{d\nu}{dx}\Big)}\nu(dx),$
(65) $\Gamma^{(\alpha)}_{ij,k}$= $T\int_{-\infty}^{\infty}\Big(\partial_{i}\partial_{j}\log{\Big(\frac{d\nu}{dx}\Big)}+\frac{1-\alpha}{2}\partial_{i}\log{\Big(\frac{d\nu}{dx}\Big)}\partial_{j}\log{\Big(\frac{d\nu}{dx}\Big)}\Big)\partial_{k}\log{\Big(\frac{d\nu}{dx}\Big)}\nu(dx).$
Proof.
For a given divergence or distance function $\mathcal{D}$, the metric tensor and the connection of information geometry [1] are obtained as follows:
(66) $g_{ij}$ $=-\mathcal{D}(\partial_{i},\tilde{\partial}_{j})|_{\tilde{\boldsymbol{\xi}}=\boldsymbol{\xi}},$
(67) $\Gamma_{ij,k}$ $=-\mathcal{D}(\partial_{i}\partial_{j},\tilde{\partial}_{k})|_{\tilde{\boldsymbol{\xi}}=\boldsymbol{\xi}},$
where $i,j,$ and $k$ run for the coordinate system $\boldsymbol{\xi}$.
As done in Theorem 1, we divide the analysis into two cases: $\sigma\neq 0$ and $\sigma=0$.
We begin with the case of $\sigma\neq 0$. By substituting Eq. (40) into Eq. (66) and Eq. (67), we obtain Eq. (60) and Eq. (61).
When the martingale condition in Eq. (5) is applied to Eq. (60) and Eq. (61), the resulting geometric structures are given by Eq. (62) and Eq. (63). The same geometric results can also be obtained by plugging Eq. (47) into Eq. (66) and Eq. (67).
In the case of $\sigma=0$, the first terms of the exponents in Eq. (40) and Eq. (47) vanish. Consequently, the first terms in Eq. (60) and Eq. (61) (or equivalently, in Eq. (62) and Eq. (63)) also become zero. A direct computation using Eq. (51) in Eq. (66) and Eq. (67) yields the same geometric structures. ∎
[…]
Corollary 2.
Let $(X_{t},\mathbb{P})_{t\in[0,T]}$ be a Lévy process on $(\Omega,\mathcal{F})$ with the Lévy triplet of $(\sigma,\nu,\gamma)$. The Fisher information matrix of Lévy processes and the $\alpha$-connection for information geometry of Lévy processes are given as
(68) $g_{ij}$ $=-\frac{4}{1-\alpha^{2}}\partial_{i}\tilde{\partial}_{j}\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})|_{\tilde{\boldsymbol{\xi}}=\boldsymbol{\xi}},$
(69) $\Gamma_{ij,k}^{(\alpha)}$ $=-\frac{4}{1-\alpha^{2}}\partial_{i}\partial_{j}\tilde{\partial}_{k}\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})|_{\tilde{\boldsymbol{\xi}}=\boldsymbol{\xi}},$
where $i,j,$ and $k$ run for the coordinate system $\boldsymbol{\xi}$.
Proof.
By substituting the $\alpha$-divergence from Corollary 2 into Eq. (66) and Eq. (67), we obtain the Fisher information matrix and the $\alpha$-connection for Lévy processes as follows:
(70) $g_{ij}$ $=-\frac{4}{1-\alpha^{2}}\Delta^{(\alpha)}_{T}(\partial_{i},\tilde{\partial}_{j}){\rm e}^{-\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})}|_{\tilde{\boldsymbol{\xi}}=\boldsymbol{\xi}},$
(71) $\Gamma_{ij,k}^{(\alpha)}$ $=-\frac{4}{1-\alpha^{2}}\Delta^{(\alpha)}_{T}(\partial_{i}\partial_{j},\tilde{\partial}_{k}){\rm e}^{-\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})}|_{\tilde{\boldsymbol{\xi}}=\boldsymbol{\xi}}.$
As mentioned earlier, when $\mathbb{P}$ and $\mathbb{Q}$ are identical, i.e., $\boldsymbol{\xi}=\tilde{\boldsymbol{\xi}}$, we have $\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})=0$. This simplifies the expressions for the Fisher information matrix and the $\alpha$-connection to
(72) $g_{ij}$ $=-\frac{4}{1-\alpha^{2}}\partial_{i}\tilde{\partial}_{j}\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})|_{\tilde{\boldsymbol{\xi}}=\boldsymbol{\xi}},$
(73) $\Gamma_{ij,k}^{(\alpha)}$ $=-\frac{4}{1-\alpha^{2}}\partial_{i}\partial_{j}\tilde{\partial}_{k}\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})|_{\tilde{\boldsymbol{\xi}}=\boldsymbol{\xi}}.$
It is worth noting that Corollary 2 remains valid even when $\alpha=\pm 1$, since the factor $\frac{4}{1-\alpha^{2}}$ cancels out during differentiation when computing the partial derivatives.
This information geometry of Lévy processes can be utilized for statistical advantages such as bias reduction [17, 26, 27] and the construction of Bayesian predictive priors [25, 35, 34, 12, 11, 29].
For instance, when the information geometry is e-flat, it is possible to employ the penalized log-likelihood approach [17]:
(74) $l^{*}(\xi)=l(\xi)+\log\mathcal{J},$
where $l(\xi)$ is log-likelihood and $\mathcal{J}$ is the Jeffreys prior.
For Lévy process geometry, the Jeffreys prior is obtained as
(75) $\mathcal{J}\propto(\det{g_{ij}})^{1/2}=\bigg(\det{\Big(-\frac{4}{1-\alpha^{2}}\Delta^{(\alpha)}_{T}(\partial_{i},\tilde{\partial}_{j})|_{\tilde{\boldsymbol{\xi}}=\boldsymbol{\xi}}\Big)}\bigg)^{1/2}.$
This Jeffreys prior can also be employed as a baseline for constructing Bayesian predictive priors. According to Komaki [25], the following Bayesian predictive priors provide improved predictive performance over the Jeffreys prior:
(76) $\tilde{\mathcal{J}}=\rho\mathcal{J},$
where $\rho$ is a superharmonic function such that $\Delta\rho<0$. The Laplace–Beltrami operator $\Delta$ is defined as
(77) $\Delta\rho=\frac{1}{(\det{g_{ij}})^{1/2}}\partial_{i}\big((\det{g_{ij}})^{1/2}g^{ij}\partial_{j}\rho\big).$
4. Examples
In this section, we present applications of the results developed in Section 3 to various Lévy processes. In particular, the examples considered are closely related to models that are widely used in finance.
4.1. Tempered stable processes
We present the geometric structures of tempered stable processes, which form a subclass of Lévy processes. Choi [8] found the information geometry of various tempered stable processes and its statistical applications in parameter estimation and Bayesian prediction. Since the examples used in Choi [8] include generalized tempered stable (GTS) process, classical tempered stable (CTS) processes and rapidly-decreasing temepered stable processes, we compare our results with those previously established in the literature.
A tempered stable distribution is characterized by a Lévy triplet of the form $(0,\nu,\gamma)$ [30], that the Lévy measure $\nu$ is represented with
(78) $\nu(dx;\boldsymbol{\xi})=t(x;\boldsymbol{\xi})\nu_{\textrm{stable}}(dx;\boldsymbol{\xi}),$
where $t(x;\boldsymbol{\xi})$ denotes the tempering function of the tempered stable process, and $\nu_{\textrm{stable}}(dx;\boldsymbol{\xi})$ represents the Lévy measure of a stable process. The Lévy measure of stable processes is given by
(79) $\nu_{\textrm{stable}}(dx;\boldsymbol{\xi})=\Big(\frac{C_{+}}{x^{a_{+}+1}}\mathbbm{1}_{x>0}(x)+\frac{C_{-}}{|x|^{a_{-}+1}}\mathbbm{1}_{x<0}(x)\Big)dx,$
where $C_{\pm}>0$, and $a_{\pm}\in(0,2)\setminus\{1\}$.
Since tempered stable processes are a special case of Lévy processes with $\sigma=0$, we can directly apply Theorem 1. From Eq. (51) in Theorem 1, the $\alpha$-divergence between two tempered stable processes is obtained as
(83) $D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=\left\{\begin{array}[]{ll}\frac{4}{1-\alpha^{2}}\Big\{1-\exp{\big[-T\int_{-\infty}^{\infty}\Big(\frac{1-\alpha}{2}\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+\frac{1+\alpha}{2}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)^{\frac{1-\alpha}{2}}\Big)\nu_{\mathbb{Q}}(dx)\big]}\Big\}&(\alpha\neq\pm 1)\\ T\int_{-\infty}^{\infty}\Big(\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)\log{\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)}-\Big(\frac{d\nu_{\mathbb{P}}}{d\nu_{\mathbb{Q}}}\Big)+1\Big)\nu_{\mathbb{Q}}(dx)&(\alpha=-1)\\ T\int_{-\infty}^{\infty}\Big(\Big(\frac{d\nu_{\mathbb{Q}}}{d\nu_{\mathbb{P}}}\Big)\log{\Big(\frac{d\nu_{\mathbb{Q}}}{d\nu_{\mathbb{P}}}\Big)}-\Big(\frac{d\nu_{\mathbb{Q}}}{d\nu_{\mathbb{P}}}\Big)+1\Big)\nu_{\mathbb{P}}(dx)&(\alpha=1)\end{array}\right.$.
Eq. (83) is consistent with the $\alpha$-divergence presented in Choi [8]. Specifically, when $\alpha=\pm 1$, the divergence coincides with the Kullback–Leibler divergence and its dual divergence found in the literature. Moreover, for $\alpha\neq\pm 1$, the $\alpha$-divergence in Choi’s work corresponds to the linear term in $T$ from Eq. (51) which is also related to Eq. (59). These expressions closely agree when $T$ is sufficiently small or when the parameters of $\mathbb{P}$ and $\mathbb{Q}$ are close enough.
Similarly, the information geometry of tempered stable processes can be derived from Theorem 2. The Fisher information matrix and the $\alpha$-connection for the information geometry of tempered stable processes are given by
(84) $g_{ij}=$ $T\int_{-\infty}^{\infty}\partial_{i}\log{\Big(\frac{d\nu}{dx}\Big)}\partial_{j}\log{\Big(\frac{d\nu}{dx}\Big)}\nu(dx),$
(85) $\Gamma^{(\alpha)}_{ij,k}$= $T\int_{-\infty}^{\infty}\Big(\partial_{i}\partial_{j}\log{\Big(\frac{d\nu}{dx}\Big)}+\frac{1-\alpha}{2}\partial_{i}\log{\Big(\frac{d\nu}{dx}\Big)}\partial_{j}\log{\Big(\frac{d\nu}{dx}\Big)}\Big)\partial_{k}\log{\Big(\frac{d\nu}{dx}\Big)}\nu(dx),$
where $i,j,$ and $k$ run for coordinate system $\boldsymbol{\xi}$. The geometry of tempered stable processes given by Eq. (84) and Eq. (85) is also identical to that in Choi [8].
As a more concrete example, let us focus on GTS processes. The GTS process uses the following tempering function [30]:
(86) $t(x;\boldsymbol{\xi})=\big({\rm e}^{-\lambda_{+}x}\mathbbm{1}_{x>0}(x)+{\rm e}^{-\lambda_{-}|x|}\mathbbm{1}_{x<0}(x)\big),$
where $\lambda_{\pm}>0$. Plugging Eq. (86) to Eq. (78) leads to the Lévy measure of GTS processes:
(87) $\nu(dx;\boldsymbol{\xi})=\Big(\frac{C_{+}{\rm e}^{-\lambda_{+}x}}{x^{a_{+}+1}}\mathbbm{1}_{x>0}(x)+\frac{C_{-}{\rm e}^{-\lambda_{-}|x|}}{|x|^{a_{-}+1}}\mathbbm{1}_{x<0}(x)\Big)dx.$
The Radon–Nikodym derivative of GTS processes is defined as
(88) $\frac{d\nu^{\mathbb{P}}}{d\nu^{\mathbb{Q}}}={\rm e}^{-(\lambda_{+}-\tilde{\lambda}_{+})x}\mathbbm{1}_{x>0}(x)+{\rm e}^{-(\lambda_{-}-\tilde{\lambda}_{-})|x|}\mathbbm{1}_{x<0}(x).$
It is noteworthy that all other parameters of two GTS processes, except for $\lambda_{\pm}$, are identical due to the existence condition of the Radon–Nikodym derivative in Proposition 1. More details on this condition can be found in Rachev et al. [30] and Choi [8].
By substituting Eq. (87) and Eq. (88) into Eq. (83), the $\alpha$-divergence between GTS processes is obtained as
(95) $D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=\left\{\begin{array}[]{ll}\frac{4}{1-\alpha^{2}}\Big\{1-\exp\big[-TC_{+}\Gamma(-a_{+})\big(\frac{1-\alpha}{2}\lambda_{+}^{a_{+}}+\frac{1+\alpha}{2}\tilde{\lambda}_{+}^{a_{+}}-(\frac{1-\alpha}{2}\lambda_{+}+\frac{1+\alpha}{2}\tilde{\lambda}_{+})^{a_{+}}\big)\\ -TC_{-}\Gamma(-a_{-})\big(\frac{1-\alpha}{2}\lambda_{-}^{a_{-}}+\frac{1+\alpha}{2}\tilde{\lambda}_{-}^{a_{-}}-(\frac{1-\alpha}{2}\lambda_{-}+\frac{1+\alpha}{2}\tilde{\lambda}_{-})^{a_{-}}\big)\big]\Big\}&(\alpha\neq\pm 1)\\ TC_{+}\Gamma(-a_{+})\big((a_{+}-1)\lambda_{+}^{a_{+}}-a_{+}\tilde{\lambda}_{+}\lambda_{+}^{a_{+}-1}+\tilde{\lambda}_{+}^{a_{+}}\big)\\ +TC_{-}\Gamma(-a_{-})\big((a_{-}-1)\lambda_{-}^{a_{-}}-a_{-}\tilde{\lambda}_{-}\lambda_{-}^{a_{-}-1}+\tilde{\lambda}_{-}^{a_{-}}\big)&(\alpha=-1)\\ TC_{+}\Gamma(-a_{+})\big((a_{+}-1)\tilde{\lambda}_{+}^{a_{+}}-a_{+}\lambda_{+}\tilde{\lambda}_{+}^{a_{+}-1}+\lambda_{+}^{a_{+}}\big)\\ +TC_{-}\Gamma(-a_{-})\big((a_{-}-1)\tilde{\lambda}_{-}^{a_{-}}-a_{-}\lambda_{-}\tilde{\lambda}_{-}^{a_{-}-1}+\lambda_{-}^{a_{-}}\big)&(\alpha=1)\end{array}\right.$.
As mentioned above, the $\alpha$-divergences for $\alpha=\pm 1$ are matched with those in Choi [8]. The $\alpha$-divergence for $\alpha\neq\pm 1$ in Choi [8] is the linear term in $T$ of Eq. (95). This is also related to Eq. (59).
Similar to the $\alpha$-divergence, the information geometry of GTS processes is found by Eq. (84) and Eq. (85). First of all, the Fisher information matrix of GTS processes is calculated as
(98) $g_{ij}=\Bigg(\begin{array}[]{cc}\frac{TC_{+}\Gamma(2-a_{+})}{\lambda_{+}^{2-a_{+}}}&0\\ 0&\frac{TC_{-}\Gamma(2-a_{-})}{\lambda_{-}^{2-a_{-}}}\end{array}\Bigg),$
where $i,j$ run for $\lambda_{+}$ and $\lambda_{-}$. Moreover, the non-vanishing components of $\alpha$-connection in GTS geometry are derived as
(99) $\Gamma^{(\alpha)}_{\lambda_{+}\lambda_{+},\lambda_{+}}$ $=-\frac{1-\alpha}{2}\frac{TC_{+}\Gamma(3-a_{+})}{\lambda_{+}^{3-a_{+}}},$
(100) $\Gamma^{(\alpha)}_{\lambda_{-}\lambda_{-},\lambda_{-}}$ $=-\frac{1-\alpha}{2}\frac{TC_{-}\Gamma(3-a_{-})}{\lambda_{-}^{3-a_{-}}}.$
The other components are vanishing. When $\alpha=0$, Eq. (99) and Eq. (100) correspond to Levi-Civita connection which is also given in Choi [8].
It is worth noting that the Fisher information matrix and the $\alpha$-connection for GTS processes coincide with those derived in Choi [8]. Moreover, the same study introduced various statistical applications of GTS geometry, including penalized likelihood estimation for bias reduction [17, 26, 27] and Bayesian predictive priors [25]. Since our geometric results are equivalent, the same conclusions apply to these statistical methodologies.
4.2. CTS processes and CGMY model
Another example discussed in Choi [8] is classical tempered stable (CTS) processes, also known in the finance literature as the CGMY model [6].
The CTS process is a special case of the GTS process, obtained by imposing the following condition:
(103) $\left\{\begin{array}[]{ll}a_{+}=a_{-}=a\\ C_{+}=C_{-}=C\end{array}\right.$.
This condition implies that both the upper and lower tails are controlled by the same tail index and scale parameter.
By applying the CTS condition from Eq. (103) to the GTS Lévy measure in Eq. (87), the Lévy measure of CTS processes is represented with
(104) $\nu(dx;\boldsymbol{\xi})=C\Big(\frac{{\rm e}^{-\lambda_{+}x}}{x^{a+1}}\mathbbm{1}_{x>0}(x)+\frac{{\rm e}^{-\lambda_{-}|x|}}{|x|^{a+1}}\mathbbm{1}_{x<0}(x)\Big)dx,$
where $C,\lambda_{+},\lambda_{-}$ are positive, and $a\in(0,2)\setminus\{1\}$.
It is straightforward to check that the Radon–Nikodym derivative of CTS processes is identical to that in Eq. (88):
(105) $\frac{d\nu^{\mathbb{P}}}{d\nu^{\mathbb{Q}}}={\rm e}^{-(\lambda_{+}-\tilde{\lambda}_{+})x}\mathbbm{1}_{x>0}(x)+{\rm e}^{-(\lambda_{-}-\tilde{\lambda}_{-})|x|}\mathbbm{1}_{x<0}(x).$
As mentioned for GTS processes, the existence of the Radon–Nikodym derivative under Proposition 1 requires that all parameters of two CTS processes, except for $\lambda_{\pm}$, are identical. Additional details regarding this condition can be found in the literature [30, 8].
The $\alpha$-divergence between CTS processes can be derived either by applying Theorem 1 or by substituting the CTS condition of Eq. (103) into Eq. (95). Leveraging the latter approach, which is also more straightforward, the $\alpha$-divergence is obtained as
(110) $D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=\left\{\begin{array}[]{ll}\frac{4}{1-\alpha^{2}}\Big\{1-\exp\big[-TC\Gamma(-a)\big(\big(\frac{1-\alpha}{2}\lambda_{+}^{a}+\frac{1+\alpha}{2}\tilde{\lambda}_{+}^{a}-(\frac{1-\alpha}{2}\lambda_{+}+\frac{1+\alpha}{2}\tilde{\lambda}_{+})^{a}\big)\\ +\big(\frac{1-\alpha}{2}\lambda_{-}^{a}+\frac{1+\alpha}{2}\tilde{\lambda}_{-}^{a}-(\frac{1-\alpha}{2}\lambda_{-}+\frac{1+\alpha}{2}\tilde{\lambda}_{-})^{a}\big)\big)\big]\Big\}&(\alpha\neq\pm 1)\\ TC\Gamma(-a)\Big(\big((a-1)\lambda_{+}^{a}-a\tilde{\lambda}_{+}\lambda_{+}^{a-1}+\tilde{\lambda}_{+}^{a}\big)+\big((a-1)\lambda_{-}^{a}-a\tilde{\lambda}_{-}\lambda_{-}^{a-1}+\tilde{\lambda}_{-}^{a}\big)\Big)&(\alpha=-1)\\ TC\Gamma(-a)\Big(\big((a-1)\tilde{\lambda}_{+}^{a}-a\lambda_{+}\tilde{\lambda}_{+}^{a-1}+\lambda_{+}^{a}\big)+\big((a-1)\tilde{\lambda}_{-}^{a}-a\lambda_{-}\tilde{\lambda}_{-}^{a-1}+\lambda_{-}^{a}\big)\Big)&(\alpha=1)\end{array}\right.$.
It is also evident that the same result can be obtained directly from Theorem 1. Moreover, similar to the case of GTS processes, the $\alpha$-divergence in Eq. (110) coincides with the results in Choi [8] when $\alpha=\pm 1$. For $\alpha\neq\pm 1$, the expression in Choi [8] corresponds to the linear term in $T$ from Eq. (110), also related to Eq. (59).
The Fisher information matrix and the $\alpha$-connection for CTS processes are also identical to those derived in Choi [8]. Specifically, plugging the CTS condition of Eq. (103) into Eq. (98) yields the information geometry of CTS processes. The Fisher information matrix of the CTS geometry is calculated as
(113) $g_{ij}=\Bigg(\begin{array}[]{cc}\frac{TC\Gamma(2-a)}{\lambda_{+}^{2-a}}&0\\ 0&\frac{TC\Gamma(2-a)}{\lambda_{-}^{2-a}}\end{array}\Bigg),$
where $i,j$ run for $\lambda_{+}$ and $\lambda_{-}$. Similarly, the non-trivial components of the $\alpha$-connection for CTS processes are given as
(114) $\Gamma^{(\alpha)}_{\lambda_{+}\lambda_{+},\lambda_{+}}$ $=-\frac{1-\alpha}{2}\frac{TC\Gamma(3-a)}{\lambda_{+}^{3-a}},$
(115) $\Gamma^{(\alpha)}_{\lambda_{-}\lambda_{-},\lambda_{-}}$ $=-\frac{1-\alpha}{2}\frac{TC\Gamma(3-a)}{\lambda_{-}^{3-a}}.$
All other components vanish. It is worth noting that the geometry derived directly from Theorem 2 coincides with the results in Eq. (113), Eq. (114), and Eq. (115).
We can apply the information geometry of CTS processes for statistical advantages.
First, since the geometry of CTS processes is e-flat, the penalized log-likelihood approach [17] can be considered as described earlier. From the Fisher information matrix in Eq. (113), the Jeffreys prior of CTS processes is represented with
(116) $\mathcal{J}(\xi)\propto\frac{TC\Gamma(2-a)}{\sqrt{(\lambda_{+}\lambda_{-})^{2-a}}}.$
Moreorever, we can find potential candidates for a superharmonic function $\rho$ from the geometry for Komaki’s Bayesian predictive priors [25]. One example is as follows:
(117) $\rho_{\pm}=\lambda_{\pm}^{k},$
where $\min(0,a-1)<k<\max(0,a-1)$. Another candidate is the positive linear combination or multiplication of the functions above:
(118) $\rho_{1}$ $=c_{1}\rho_{+}+c_{2}\rho_{-},$
(119) $\rho_{2}$ $=\rho_{+}\rho_{-},$
where $c_{+}>0$ and $c_{-}>0$.
4.3. Variance gamma processes
Variance gamma (VG) processes [28] are also widely used in financial modeling. The VG process is a Lévy process characterized by the triplet $(0,\nu,\gamma)$. However, it does not fall within the class of tempered stable processes.
The Lévy measure of variance gamma processes is represented with
(120) $\nu(dx;\boldsymbol{\xi})=C\Big(\frac{{\rm e}^{-\lambda_{+}x}}{x}\mathbbm{1}_{x>0}(x)+\frac{{\rm e}^{-\lambda_{-}|x|}}{|x|}\mathbbm{1}_{x<0}(x)\Big)dx,$
where $C,\lambda_{+},\lambda_{-}$ are positive.
However, the integral of Eq. (120) over $(-\infty,\infty)$ diverges making it ill-defined for computing the $\alpha$-divergence directly. To regularize the expression, we introduce a small positive parameter $a$ and consider the modified Lévy measure:
(121) $\nu(dx;\boldsymbol{\xi})=C\Big(\frac{{\rm e}^{-\lambda_{+}x}}{x^{1+a}}\mathbbm{1}_{x>0}(x)+\frac{{\rm e}^{-\lambda_{-}|x|}}{|x|^{1+a}}\mathbbm{1}_{x<0}(x)\Big)dx,$
This modified measure is integrable over the real line, allowing the use of standard divergence calculations, and the original measure can be recovered by taking the limit $a\to 0$.
It is noteworthy that Eq. (121) coincides with the Lévy measure of the CTS distribution, Eq. (104). This structural similarity suggests that one could formally obtain the VG process as a limiting case of the CTS process by setting $a=0$. While this makes it tempting to directly substitute $a=0$ into the $\alpha$-divergence expression for CTS processes, Eq. (110), caution is necessary as the Gamma function is undefined at zero. Additionally, $a=0$ is not allowed for CTS and GTS processes.
To address the behavior of the Gamma function near $a=0$, we take the following approximation:
(124) $\left\{\begin{array}[]{ll}\Gamma(a)&\approx\frac{1}{a}-\gamma+\mathcal{O}(a)\\ \lambda^{a}&\approx 1+a\log{\lambda}+\mathcal{O}(a^{2})\end{array}\right$.,
where $\gamma$ is the Euler–Mascheroni constant.
By applying Eq. (124) to Eq. (110) and taking the limit as $a\to 0$, we obtain the $\alpha$-divergence between two variance gamma processes as:
(129) $D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=\left\{\begin{array}[]{ll}\frac{4}{1-\alpha^{2}}\Big\{1-\exp\big[-TC\big(\big(-\frac{1-\alpha}{2}\log{\lambda_{+}}-\frac{1+\alpha}{2}\log{\tilde{\lambda}}+\log{(\frac{1-\alpha}{2}\lambda_{+}+\frac{1+\alpha}{2}\tilde{\lambda}_{+})\big)}\\ +\big(-\frac{1-\alpha}{2}\log\lambda_{-}-\frac{1+\alpha}{2}\log\tilde{\lambda}_{-}+\log(\frac{1-\alpha}{2}\lambda_{-}+\frac{1+\alpha}{2}\tilde{\lambda}_{-})\big)\big)\big]\Big\}&(\alpha\neq\pm 1)\\ TC\Big(\big(\tilde{\lambda}_{+}/\lambda_{+}-1+\log{\lambda_{+}}-\log{\tilde{\lambda}_{+}}\big)+\big(\tilde{\lambda}_{-}/\lambda_{-}-1+\log{\lambda_{-}}-\log{\tilde{\lambda}_{-}}\big)\Big)&(\alpha=-1)\\ TC\Big(\big(\lambda_{+}/\tilde{\lambda}_{+}-1+\log{\tilde{\lambda}_{+}}-\log{\lambda_{+}}\big)+\big(\lambda_{-}/\tilde{\lambda}_{-}-1+\log{\tilde{\lambda}_{-}}-\log{\lambda_{-}}\big)\Big)&(\alpha=1)\end{array}\right.$.
By taking the limit of $a\to 0$ to Eq. (113), the Fisher information matrix of variance gamma processes is represented with
(132) $g_{ij}=\Bigg(\begin{array}[]{cc}\frac{TC}{\lambda_{+}^{2}}&0\\ 0&\frac{TC}{\lambda_{-}^{2}}\end{array}\Bigg),$
where $i,j,$ and $k$ run for $\lambda_{\pm}$. The $\alpha$-connection of variance gamma process geometry is given as
(133) $\Gamma^{(\alpha)}_{\lambda_{+}\lambda_{+},\lambda_{+}}$ $=-(1-\alpha)\frac{TC}{\lambda_{+}^{3}},$
(134) $\Gamma^{(\alpha)}_{\lambda_{-}\lambda_{-},\lambda_{-}}$ $=-(1-\alpha)\frac{TC}{\lambda_{-}^{3}}.$
All other components are vanishing.
As with GTS and CTS processes, the information geometry of variance gamma processes enables a range of statistical applications. Since the VG geometry is also e-flat, we can consider the penalized likelihood method for bias reduction [17]. Based on the Fisher information matrix for variance gamma processes given by Eq. (132), the corresponding Jeffreys prior is obtained as
(135) $\mathcal{J}(\xi)\propto\frac{TC}{\lambda_{+}\lambda_{-}},$
and it can be applied to Eq. (74) for bias reduction.
This Jeffreys prior can be leveraged to construct Bayesian predictive priors, as discussed in Komaki [25]. Moreover, superharmonic functions for variance gamma processes can be obtained by taking the limit $a\to 0$ in the corresponding superharmonic functions for CTS processes:
(136) $\rho_{\pm}=\lambda_{\pm}^{k},$
where 1<k<0-1<k<0. Similarly, the positive linear combination or multiplication of the functions above can be considered:
(137) $\rho_{1}$ $=c_{1}\rho_{+}+c_{2}\rho_{-},$
(138) $\rho_{2}$ $=\rho_{+}\rho_{-},$
where $c_{+}>0$ and $c_{-}>0$.
5. Conclusion
In this paper, we have developed an information-geometric framework for Lévy processes, significantly extending precedent works focused on the information geometry of tempered stable processes [8] and the Kullback–Leibler divergence of Lévy processes [14]. By deriving the $\alpha$-divergence for Lévy processes, we constructed the corresponding Fisher information matrices and $\alpha$-connections, uncovering the underlying differential-geometric structure of this broad and important class of stochastic processes. Statistical advantages from obtaining the information geometry were also discussed.
Moreover, we demonstrated the utility of this geometric framework by applying it to several prominent Lévy models widely used in finance, including tempered stable processes, CGMY models, and variance gamma processes. For each of these models, we characterized their information geometry and discussed potential applications to bias-reduced estimation and Bayesian predictive priors that are tools relevant to model calibration, estimation robustness, and risk measurement in financial contexts. These examples underscore how information geometry functions not only as a theoretical framework but also as a practical methodology in quantitative modeling.
In closing, this work contributes to the growing interdisciplinary study of stochastic process theory, information geometry, and applied probability. By embedding Lévy processes within the information-geometric framework, we offer a novel foundation for analyzing and comparing complex models exhibiting jumps and non-Gaussian behavior. We hope that this study inspires further research at the intersection of geometry, statistics, and stochastic modeling.
Acknowledgment
We thank Hyangju Kim and Young Shin Kim for useful discussions on Lévy processes.